# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

### v2

@@ Abstract
In legal decisions, split votes (SV) occur when judges cannot reach a unanimous decision, posing a difficulty for lawyers who must navigate diverse legal arguments and opinions. In high-stakes domains, understanding the alignment of perceived difficulty between humans and AI systems is crucial to build trust. However, existing NLP calibration methods focus on a classifier’s awareness of predictive performance, measured against the human majority class, overlooking inherent human label variation (HLV). This paper explores split votes as naturally observable human disagreement and value pluralism. We collect judges’ vote distributions from the European Court of Human Rights (ECHR), and present SV-ECHR [^1] Our dataset and code will be available upon publication. a case outcome classification (COC) dataset with SV information. We build a taxonomy of disagreement with SV-specific subcategories. We further assess the alignment of perceived difficulty between models and humans, as well as confidence- and human-calibration of COC models. We observe limited alignment with the judge vote distribution. To our knowledge, this is the first systematic exploration of calibration to human judgements in legal NLP. Our study underscores the necessity for further research on measuring and enhancing model calibration considering HLV in legal decision tasks.
@@
[…]
@@ 1 Introduction
The task of Case Outcome Classification (COC) involves the classification of legal case outcomes based on textual descriptions of their facts. While achieving high performance is desirable, in high-stakes domains such as legal and medical decisions, the quantification of a model’s predictive confidence, or conversely, its uncertainty, is particularly valuable. The incorporation of uncertainty estimation allows experts to make more informed decisions, especially when the model may be uncertain or where the consequences of a misdiagnosis are significant. Evaluating whether models are aware of their limitations is known as calibration Guo et al. (2017); Desai and Durrett (2020). It assesses the extent to which predictive probabilities accurately reflect the likelihood of a prediction being correct. Models can opt to abstain when the uncertainty exceeds a predefined threshold — a method commonly referred to as selective classification El-Yaniv et al. (2010); Geifman and El-Yaniv (2017).
Current NLP research focuses on prediction confidence and calibration to assess a classifier’s awareness of its predictive performance only. This evaluation is commonly conducted against the human majority class. However, recent developments in NLP research have shed light on the prevalence of inherent human label variation (HLV) Plank (2022), observing disagreement [^2] The term HLV embraces disagreement and plausible variation. In this paper, we use these terms interchangeably. across various tasks Uma et al. (2021). Scholars in the field argue for the acknowledgment and acceptance of HLV, as it mirrors the diverse and pluralistic nature of human values Sorensen et al. (2024). Notably, Baan et al. 2022 has demonstrated that widely-used calibration metrics may not be suitable for datasets characterized by inherent human disagreement.
@@ 2 Related Work
COC has been referred to as Legal Judgment Prediction (LJP) in previous research Medvedeva et al. (2020); T.y.s.s et al. (2023). There exist numerous works involving corpora from ECtHR Aletras et al. (2016); Chalkidis et al. (2019); Tyss et al. (2023). All of these approaches mainly focus on the COC performance of models, which is commonly measured against the human majority class. Our SV-ECtHR dataset extends this work and contains the nuanced judges’ vote information. We proceed to use it to systematically investigate disagreement among judges and the alignment of perceived difficulty between models and judges.
@@
Disagreement / Human Label Variation is receiving growing attention in mainstream NLP. Various works highlight the presence of HLV, emphasizing the abundance and plausibility of such human disagreements Pavlick and Kwiatkowski (2019); Plank (2022). Researchers advocate for embracing HLV for more trustworthy AI Talat et al. (2022); Casper et al. (2023). As real-world applications are used to assist diverse audiences, it becomes crucial to investigate and include pluralistic human values in NLP systems Sorensen et al. (2023). This motivates our study of disagreements among judges’ split-votes in the legal decision process. Recently, several studies have proposed task-specific taxonomies to identify potential sources of disagreement in various NLP tasks Uma et al. (2021); Sandri et al. (2023); Jiang and Marneffe (2022). Building on the meta-analysis of these existing taxonomies, Xu et al. 2023b generalize them to two layers of task-agnostic categories and introduce task-specific categories for COC rationale annotation. In this work, we study judges’ disagreements in case decision votes, presenting a taxonomy of disagreement with SV-specific subcategories.
@@
Perceived Difficulty Regarding model difficulty, there is increasing interest in identifying difficult data instances. Various techniques, such as Influence Functions Koh and Liang (2017) and training loss Han et al. (2018) have been proposed to identify the difficulty of data instances to a certain model. Pointwise V-usable information (PVI) is a recently introduced difficulty metric Ethayarajh et al. (2022), which incorporates mutual information and other types of informativenessXu et al. (2020). Despite its recent introduction, PVI has gained significant attention and proven effective in various tasks, including rationale evaluation Prasad et al. (2023) and data selection for augmentation Lin et al. (2023). In this work, we leverage PVI to evaluate the alignment of perceived difficulty between human and COC models.
@@
Calibration Existing calibration studies primarily focus on a classifier’s awareness of its predictive performance, commonly evaluated against the human majority class Guo et al. (2017); Desai and Durrett (2020). Baan et al. 2022 argue that calibrating against the human majority class is not meaningful in settings with inherent HLV. In our work, we assess the performance and human calibration of COC models in the context of split-vote, a naturally occurring HLV. To the best of our knowledge, this is the first systematic exploration of human calibration in the legal domain.
@@ 3 Legal and Linguistic Backgrounds
Sources of Disagreement in Legal Decisions Recent research has delved into sources of HLV and disagreement across various NLP tasks from a linguistic perspective. The classic framework “Triangle of Reference" Ogden and Richards (1923) has been widely adopted to categorize disagreement sources in classification tasks Aroyo and Welty (2015); Jiang and Marneffe (2022). This concept primarily addresses the relationship between linguistic symbols and the corresponding objects they represent. To study uncertainty sources in NLG tasks, Baan et al. 2023 extend it to the “Double Triangle of Language Production” catering to the complexities of language generation. However, the legal decision-making process introduces an additional layer of complexity. Legal scholars conceptualize the decision-making process in case-law as circular Ichim (2019). To better capture this nature, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4 in App A), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the continuous reproduction of the main normative model over time.
@@ 4 The SV-ECHR Dataset
Dataset Collection We extract the judge votes distribution for each alleged article [^3] Following Chalkidis et al. 2022a; Santosh et al. 2022 We use only the 10 most prominent ECHR articles. from the public database HUDOC [^4] https://hudoc.echr.coe.int using regular expressions. The information about judge votes is always present in the conclusion section and generally follows certain patterns (See more details in App B). We did a two-round quality assessment. [^5] App C offers details on the quality assessment process. The F1 score of our regular expression rules increased from 0.81 in the first round to 0.98 in the second round.
Dataset Analysis We then augment the ECHR A|B dataset Santosh et al. (2022) with our collected vote information according to the document ID. We name the augmented dataset SV-ECHR. It consists of 11k case fact descriptions along with target label information about which convention articles have been alleged to be violated (task B), and which the court has eventually found to be violated (task A), and the judges vote distribution of each alleged article. The dataset is chronologically split into training (2001–2016), validation (2016–2017) and test set (2017-2019) with 9k, 1k, 1k cases respectively. The label set includes 10 prominent ECHR convention articles. On average, each case has around 1.6 alleged articles. Among all 17,604 alleged case-alleged article pairs (hereafter pair), only round 7% are split-voted. These numbers underscore the significant label imbalance within the dataset. Additional statistics can be found in Tab 4 in App D. To investigate the extent of judges’ disagreement in SV-ECHR, we assess the entropies of vote distributions for each pair. For the detailed calculation and histogram of the entropy distribution, see App F. We see a large share ($\sim$ 60% of chamber votes) of single dissenting votes, where only one single judge voted differently than the six judge majority. A similar pattern is observed for Grand Chamber cases involving 17 judges. This pattern aligns with other observations made in empirical legal scholarship (Fobbe 2022, see App G).
Correction of Inconsistent Metadata Xu et al. 2023b already pointed out inconsistent allegation information in HUDOC. During our quality control process, we found further inconsistent violation information in approximately 2% of the training set. We updated SV-ECHR with the correct metadata. [^6] See App E for more details of our metadata correction Our finding of such inconsistencies calls for mindful data curation when developing COC datasets.
@@ 5.1 Disagreement Taxonomy
We utilize the disagreement taxonomy from Xu et al. 2023b to analyze the reasons behind judges’ split votes. Fig 2 displays the taxonomy, with two adapted task-agnostic levels and our expanded split-vote specific subcategories and proxy variables. [^7] For a comprehensive understanding of each taxonomy category, we direct the reader to Xu et al. 2023b In the following sections, we provide detailed explanations for each SV-specific category.
@@ 5.1.1 The Data
Genuine Ambiguity is attributed to Normative Uncertainty of case law in the context of legal NLP (Xu et al. 2023b), which emerges when the court is presented with the possibility of justifying an outcome through multiple legal source interpretations and argumentation. Its occurrence is not uncommon in ECtHR judgments due to the deliberate drafting of the convention in a ‘flexible’ manner to allow tailoring the interpretation to domestic specificities based on the subsidiarity and margin of appreciation principles. It should be noted that there exists an analog factual ambiguity where the facts of the case are unclear based on limited (or contradicting) evidence provided. As the ECtHR does not engage in evidentiary reasoning, this aspect is out of scope for this work.
Text Instantiation covers inconsistency, incompleteness, or biases during text production (i.e. judgment document drafting). When reviewing case files, judges analyze materials that encapsulate the factual background, legal arguments, and evidence presented by both parties. However, these documents are usually produced by the Court Registry, priming the language accordingly. The likelihood of encountering textual unclarity, such as inconsistent framing of facts, increases with the length and complexity of the docket. We refer to this challenge as Narrative Complexity.
Context Sensitivity refers to the Specific Legal Context in the legal domain, characterized by undefined and controversial social-political factors for which judges may not find a precisely tailored legal explanation. We use the term ‘Context Specificity’ because pinpointing the exact arrangement of factors (i.e., attitudinal, normative, and strategic) influencing context in a specific decision is challenging Shapiro (2002). Certain social and political factors may only become contentious in a particular context and arrangement of factors.
@@ 5.1.2 The Annotator/Judge
Disagreements can also arise due to variation in human behavior, which we systematize as Noise, Subjectivity and Longitudinal behavioural change. Noise covers errors due to annotator’s Sloppy Annotation or Interface Limitation Uma et al. (2021); Sandri et al. (2023). We expect negligence-related noise and interface limitations to be negligible in naturally occurring ECtHR judge votes. [^8] Empirical legal scholarship has been investigating the relationship between judge productivity and caseload, including its ramifications on decision quality (e.g., Engel and Weinshall 2020). To the best of our knowledge, no such evidence has been documented in the ECtHR context.
Subjectivity The ECtHR is composed of judges from the 46 member states with diverse legal traditions and cultural backgrounds. Previous work in general NLP demonstrated how annotators’ demographic identities can influence how they label toxicity in text Sap et al. (2022). The judges’ Cultural Background, stemming from their legal training background and political consideration from their native countries may lead to variations in their rulings compared to those of their colleagues Voeten (2008). [^9] See discussion about National Judge in App I Divergent opinions may arise when, for example, two judges from the same country prioritizing different values. Pluralistic Human Values have recently been made a primary object of mainstream NLP research Sorensen et al. (2024). Political science research has explored this in the context of quantitative justice ideal point estimation on the US Supreme Court (SCOTUS, typically a scalar dimension between liberal and conservative, see Segal and Cover 1989, Martin and Quinn 2002). More similar to our focus on agreement, Ruger et al. 2004 trained an ensemble of court-level and judge-specific decision trees, and observed different performance on conservative and liberal judges.
@@
The Longitudinal Dimension in Xu et al. 2023b considers two aspects: Individual Behavioral Change over time and Case Law Change. The latter involves shifts in the collective societal attitude towards specific phenomena over time, including changes in laws, policies, and cultural norms. In the legal domain, legal professionals must continually adapt their knowledge, strategies and reasoning to align with the ongoing evolvement of jurisprudence and applicable statutory law. Due to the limited scope of this work, we defer a comprehensive longitudinal study to future research.

### v3

@@ Abstract
In legal decisions, split votes (SV) occur when judges cannot reach a unanimous decision, posing a difficulty for lawyers who must navigate diverse legal arguments and opinions. In high-stakes domains, understanding the alignment of perceived difficulty between humans and AI systems is crucial to build trust. However, existing NLP calibration methods focus on a classifier’s awareness of predictive performance, measured against the human majority class, overlooking inherent human label variation (HLV). This paper explores split votes as naturally observable human disagreement and value pluralism. We collect judges’ vote distributions from the European Court of Human Rights (ECHR), and present SV-ECHR [^1] Our dataset and code is available at https://github.com/TUMLegalTech/SplitVote_ECHR a case outcome classification (COC) dataset with SV information. We build a taxonomy of disagreement with SV-specific subcategories. We further assess the alignment of perceived difficulty between models and humans, as well as confidence- and human-calibration of COC models. We observe limited alignment with the judge vote distribution. To our knowledge, this is the first systematic exploration of calibration to human judgements in legal NLP. Our study underscores the necessity for further research on measuring and enhancing model calibration considering HLV in legal decision tasks.
@@
Shanshan Xu1,2∗, Santosh T.Y.S.S1,Oana Ichim3 Barbara Plank4,5, Matthias Grabmair1 1Technical University of Munich, Germany 2ELTEMATE 3Graduate Institute of International and Development Studies, Switzerland 4IT University of Copenhagen, Denmark 5LMU Munich & Munich Center for Machine Learning (MCML), Germany
**footnotetext: Work done during internship at ELTEMATE
@@ 1 Introduction
The task of Case Outcome Classification (COC) involves the classification of legal case outcomes based on textual descriptions of their facts. While achieving high performance is desirable, in high-stakes domains such as legal and medical decisions, the quantification of a model’s predictive confidence, or conversely, its uncertainty, is particularly valuable. It allows experts to make more informed decisions, especially when the model may be uncertain or where the consequences of a misdiagnosis are significant. Evaluating whether models are aware of their limitations is known as assessing their uncertainty, with popular methods such as calibration Guo et al. (2017); Desai and Durrett (2020). Calibration assesses the extent to which predictive probabilities accurately reflect the likelihood of a prediction being correct. Models can opt to abstain when the uncertainty exceeds a predefined threshold — a method commonly referred to as selective classification El-Yaniv et al. (2010); Geifman and El-Yaniv (2017).
Current NLP research focuses on prediction confidence and calibration to assess a classifier’s awareness of its predictive performance only. This evaluation is commonly conducted against the human majority class. However, recent developments in NLP research have shed light on the prevalence of inherent human label variation (HLV) Plank (2022), observing disagreement***The term HLV embraces disagreement and plausible variation. In this paper, we use these terms interchangeably. across various tasks Uma et al. (2021). Scholars in the field argue for the acknowledgment and acceptance of HLV, as it mirrors the diverse and pluralistic nature of human values Sorensen et al. (2024). Notably, Baan et al. 2022 has demonstrated that widely-used calibration metrics may not be suitable for datasets characterized by inherent human disagreement.
@@ 2 Related Work
has been referred to as Legal Judgment Prediction (LJP) in previous research Medvedeva et al. (2020); T.y.s.s et al. (2023a). There exist numerous works involving corpora from ECtHR Aletras et al. (2016); Chalkidis et al. (2019); T.y.s.s et al. (2023b). All of these approaches mainly focus on the COC performance of models, which is commonly measured against the human majority class. Our SV-ECtHR dataset extends this work and contains the nuanced judges’ vote information. We proceed to use it to systematically investigate disagreement among judges and the alignment of perceived difficulty between models and judges.
Disagreement / Human Label Variation
@@
is receiving growing attention in mainstream NLP. Various works highlight the presence of HLV, emphasizing the abundance and plausibility of such human disagreements Pavlick and Kwiatkowski (2019); Plank (2022). Researchers advocate for embracing HLV for more trustworthy AI Talat et al. (2022); Casper et al. (2023). As real-world applications are used to assist diverse audiences, it becomes crucial to investigate and include pluralistic human values in NLP systems Sorensen et al. (2023). This motivates our study of disagreements among judges’ split-votes in the legal decision process. Recently, several studies have proposed task-specific taxonomies to identify potential sources of disagreement in various NLP tasks Uma et al. (2021); Sandri et al. (2023); Jiang and Marneffe (2022). Building on the meta-analysis of these existing taxonomies, Xu et al. 2023b generalize them to two layers of task-agnostic categories and introduce task-specific categories for COC rationale annotation. In this work, we study judges’ disagreements in case decision votes, presenting a taxonomy of disagreement with SV-specific subcategories.
Perceived Difficulty
@@
Regarding model difficulty, there is increasing interest in identifying difficult data instances. Various techniques, such as Influence Functions Koh and Liang (2017) and training loss Han et al. (2018) have been proposed to identify the difficulty of data instances to a certain model. Pointwise V-usable information (PVI) is a recently introduced difficulty metric Ethayarajh et al. (2022), which incorporates mutual information and other types of informativenessXu et al. (2020). Despite its recent introduction, PVI has gained significant attention and proven effective in various tasks, including rationale evaluation Prasad et al. (2023) and data selection for augmentation Lin et al. (2023). In this work, we leverage PVI to evaluate the alignment of perceived difficulty between human and COC models.
Calibration
@@
Recently in mainstream NLP, researchers posit that the overall reliability of a model is determined by two sides: 1) trustworthiness, which is addressed through models’ confidence measuring, and 2) fairness, which is addressed through their confidence alignment with humans Baan et al. (2024). Existing calibration studies primarily focus on a classifier’s confidence of its predictive performance, commonly evaluated against the human majority class Guo et al. (2017); Desai and Durrett (2020). Baan et al. 2022 instead argue that calibrating against the human majority class is not meaningful in settings with inherent HLV. In our work, we assess the performance and human calibration of COC models in the context of split-vote, a naturally occurring HLV. To the best of our knowledge, this is the first systematic exploration of human calibration in the legal domain.
Uncertainty Evaluation
Within the deep learning community, uncertainty is often classified into two types: aleatoric and epistemic uncertainty. Epistemic (or model) uncertainty arises from a lack of knowledge about the best model, often exacerbated by out-of-distribution examples. Aleatoric uncertainty, on the other hand, stems from inherent ambiguity and can be considered as the variability in experiment outcomes Houlsby et al. (2011); Gal et al. (2016). As SV is due to inherent disagreement among judges, we regard it as aleatoric uncertainty. Aleatoric uncertainty is typically quantified through metrics such as Entropy Gal et al. (2016) and Softmax Response Geifman and El-Yaniv (2017), as recognized in prior works Malinin and Gales (2018). However, studies show that these methods are based on predictive entropy, and actually measure total uncertainty, combining both epistemic and aleatoric uncertainty Van Amersfoort et al. (2020). Only when we have prior knowledge that either aleatoric or epistemic uncertainty is low, can we use predictive entropy as a suitable measure for the other type Mukhoti et al. (2023). Given our small, label-imbalanced, and temporally shifted dataset, we choose to refrain from utilizing these methods to directly measure the aleatoric uncertainty of our cases. Instead, we opt to assess difficulty using PVI, which measures predictive entropy and can be considered an equivalent for total uncertainty.
@@ 3 Legal and Linguistic Backgrounds
Sources of Disagreement in Legal Decisions
Recent research has delved into sources of HLV and disagreement across various NLP tasks from a linguistic perspective. The classic framework “Triangle of Reference” Ogden and Richards (1923) is widely adopted to categorize disagreement sources in classification tasks Aroyo and Welty (2015); Jiang and Marneffe (2022). This concept primarily addresses the relationship between linguistic symbols and the corresponding objects they represent. To study uncertainty sources in NLG tasks, Baan et al. 2023 extend it to the “Double Triangle of Language Production” catering to the complexities of language generation. However, the legal decision-making process introduces an additional layer of complexity. Legal scholars conceptualize the decision-making process in case-law as circular Ichim (2019). To better capture this nature, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 2), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the continuous reproduction of the main normative model over time.
@@ 4 The SV-ECHR Dataset
Dataset Collection
We extract the judge votes distribution for each alleged article***Following Chalkidis et al. 2022a; Santosh et al. 2022 We use only the 10 most prominent ECHR articles. from the public database HUDOC***https://hudoc.echr.coe.int using regular expressions. The information about judge votes is always present in the conclusion section and generally follows certain patterns (See more details in App A). We did a two-round quality assessment.***App B offers details on the quality assessment process. The F1 score of our regular expression rules increased from 0.81 in the first round to 0.98 in the second round.
Dataset Analysis
We then augment the ECHR A|B dataset Santosh et al. (2022) with our collected vote information according to the document ID. We name the augmented dataset SV-ECHR. It consists of 11k case fact descriptions along with target label information about which convention articles have been alleged to be violated (task B), and which the court has eventually found to be violated (task A), and the judges vote distribution of each alleged article. The dataset is chronologically split into training (2001–2016), validation (2016–2017) and test set (2017-2019) with 9k, 1k, 1k cases respectively. The label set includes 10 prominent ECHR convention articles. On average, each case has around 1.6 alleged articles. Among all 17,604 alleged case-alleged article pairs (hereafter pair), only around 7% are split-voted. These numbers underscore the significant label imbalance within the dataset. Additional statistics can be found in Tab 4 in App C. To investigate the extent of judges’ disagreement in SV-ECHR, we assess the entropies of vote distributions for each pair. For the detailed calculation and histogram of the entropy distribution, see App E. We see a large share ($\sim$ 60% of chamber votes) of single dissenting votes, where only one single judge voted differently than the six judge majority. A similar pattern is observed for Grand Chamber cases involving 17 judges. This pattern aligns with other observations made in empirical legal scholarship (Fobbe 2022, see App F).
Correction of Inconsistent Metadata
Xu et al. 2023b already pointed out inconsistent allegation information in HUDOC. During our quality control process, we found further inconsistent violation information in approximately 2% of the training set. We updated SV-ECHR with the correct metadata.***See App D for more details of our metadata correction. Our finding of such inconsistencies calls for mindful data curation when developing COC datasets.
@@ 5.1 Disagreement Taxonomy
We use the disagreement taxonomy from Xu et al. 2023b to analyze the reasons behind judges’ split votes. Fig 3 displays the taxonomy, with two adapted task-agnostic levels and our expanded split-vote specific subcategories and proxy variables.***For a comprehensive understanding of each taxonomy category, we direct the reader to Xu et al. 2023b In the following sections, we provide detailed explanations for each SV-specific category.
@@ 5.1.1 The Data
Genuine Ambiguity
is attributed to Normative Uncertainty of case law in the context of legal NLP (Xu et al. 2023b), which emerges when the court is presented with the possibility of justifying an outcome through multiple legal source interpretations and argumentation. Its occurrence is not uncommon in ECtHR judgments due to the deliberate drafting of the convention in a ‘flexible’ manner to allow tailoring the interpretation to domestic specificities based on the subsidiarity and margin of appreciation principles. It should be noted that there exists an analog factual uncertainty where the facts of the case are unclear based on limited (or contradicting) evidence provided. As the ECtHR does not engage in evidentiary reasoning, this aspect is out of scope for this work.
Text Instantiation
covers inconsistency, incompleteness, or biases during text production (i.e. judgment document drafting). When reviewing case files, judges analyze materials that encapsulate the factual background, legal arguments, and evidence presented by both parties. However, these documents are usually produced by the Court Registry, priming the language accordingly. The likelihood of encountering textual unclarity, such as inconsistent framing of facts, increases with the length and complexity of the docket. We refer to this challenge as Narrative Complexity.
Context Sensitivity
refers to the Specific Legal Context, characterized by undefined and controversial social-political factors for which judges may not find a precisely tailored legal explanation. We use the term ‘Context Specificity’ because pinpointing the exact arrangement of factors (i.e., attitudinal, normative, and strategic) influencing context in a specific decision is challenging Shapiro (2002). Certain social and political factors may only become contentious in a particular context and arrangement of factors.
@@ 5.1.2 The Annotator/Judge
Disagreements can also arise due to variation in human behavior, which we systematize as Noise, Subjectivity and Longitudinal behavioural change. Noise covers errors due to annotator’s Sloppy Annotation or Interface Limitation Uma et al. (2021); Sandri et al. (2023). We expect negligence-related noise and interface limitations to be insignificant in naturally occurring ECtHR judge votes.***Empirical legal scholarship has been investigating the relationship between judge productivity and caseload, including its ramifications on decision quality (e.g., Engel and Weinshall 2020). To the best of our knowledge, no such evidence has been documented in the ECtHR context.
Subjectivity
The ECtHR is composed of judges from the 46 member states with diverse legal traditions and cultural backgrounds. Previous work in general NLP demonstrated how annotators’ demographic identities can influence how they label toxicity in text Sap et al. (2022). The judges’ Cultural Background, stemming from their legal training and political consideration from their native countries may lead to variations in their rulings compared to those of their colleagues Voeten (2008).***See discussion about National Judge in App H Divergent opinions may arise when, for example, two judges from the same country prioritizing different values. Pluralistic Human Values have recently been made a primary object of mainstream NLP research Sorensen et al. (2024). Political science research has explored this in the context of quantitative justice ideal point estimation on the US Supreme Court (SCOTUS, typically a scalar dimension between liberal and conservative, see Segal and Cover 1989, Martin and Quinn 2002). More similar to our focus on agreement, Ruger et al. 2004 trained an ensemble of court-level and judge-specific decision trees, and observed different performance on conservative and liberal judges.
@@
The Longitudinal Dimension
in Xu et al. 2023b considers two aspects: Individual Behavioral Change over time and Case Law Change. The latter involves shifts in the collective societal attitude towards specific phenomena over time, including changes in laws, policies, and cultural norms. In the legal domain, legal professionals must continually adapt their knowledge, strategies and reasoning to align with the ongoing evolvement of jurisprudence and applicable statutory law. Due to the limited scope of this work, we defer a comprehensive longitudinal study to future research.